2. The half-life of aspirin in a person's bloodstream is about 15 minutes. A person takes a 325 milligram Bayer Aspirin.
a. Find an equation to represent this situation. Let f(t) be the amount of aspirin (in milligrams) in the bloodstream if t is the number of hours since it entered the bloodstream.
b. When will there be no aspirin left in the person’s bloodstream? (Show your algebra.)

c. What is the domain and range for this situation?



Answer by stanbon(57214)

The half-life of aspirin in a person's bloodstream is about 15 minutes. A person takes a 325 milligram Bayer Aspirin.
a. Find an equation to represent this situation. Let f(t) be the amount of aspirin (in milligrams) in the bloodstream if t is the number of hours since it entered the bloodstream.
---
f(t) = f(0)(1/2)^(t/(1/4))
--------------------------------
b. When will there be no aspirin left in the person’s bloodstream? (Show your algebra.)
----
325(1/2)^(4t) = 0
(1/2)^(4t) = 0
---
There is no solution as no power of (1/2) is equal to zero.
---
c. What is the domain and range for this situation?
Domain : t >= 0
Range: 0 < f(t) <= 325
============================
Cheers,
Stan H.